Chapter 18, Problem 05 Cosider three identical metal spheres A, B, and C. Sphere A carries a charge of +Sq. Sphere B cames a charge of g, sphere C carries no net charge. Soteres A and B·е touched together and then separated. Sphere C is then touched to sphere A and separated from it. Lastly, sphere C is touched to sphere B and separated from it. (a) What is the ratio of the final charge on sphere C to q? What is the ratio of the final total charge on the three spheres to q (b) before they are allowed to touch each other and (e) after they have touched? Units Units Units (a) Number (b) Number (c) Number Click if you would like to Show Work for this question: Ooen Show Work
Answer #1

Whenever a set of metal spheres touch each other, the total of their charges will equally shared.

Now,

When A and B are in contact, their total charge becomes (+5q) + (-1q) = 4q, which split between themselves each have +2q charge.

Now then C and A are in contact, again their total charge becomes (+2q) + 0q = + 2q, which split between them , each have +1q charge.

In the end, C and B are in contact , So again total charge is (+1q) + (+2q) = +3q , so each have

+1.5 q.

So the charge on C is +1.5 q. (ans of a)

Part B)

The final total charge before touched is

= +5q - 1q + 0q

= +4q

Part C)

The final total charge after touched is

= +1q + 1.5 q + 1.5q

= +4q